OpenStudy (anonymous):

HELLPPP PLEASEEEEE! MEDAL<3Let g(x) be the reflection of f(x) = x2 +3 in the x-axis. What is the function rule for g(x)? A. g(x) = -x2 - 3 B. g(x) = -x2 + 3 C. g(x) = x2 - 3 D. g(x) = x2 + 3

OpenStudy (anonymous):

when flipping over the x axis, the parabola would be open to the bottom when the variable ^2 is negative, so when the x2 is opposite of the original, the parabola would be a reflection. B shows this

OpenStudy (johnweldon1993):

When flipping over the x-axis we follow the rule \[\large f(x) \rightarrow -f(x)\] So for your equation \[\large f(x) = x^2 + 3\] \[\large -f(x) = x^2 + 3 \rightarrow -x^2 - 3 \]

OpenStudy (johnweldon1993):

So I receive the answer A